### v1

@@ 5 Quasi-geodesic based retraction on $\mathrm{iSt}(k,n)$
Let us first formally define a retraction on $\mathrm{iSt}(k,n)$ AbsiMS08 ; AdleDMMS02 . It is a smooth mapping $\mathcal{R}$ from the tangent bundle $T\mathrm{iSt}(k,n)\coloneqq\bigcup_{X\in\mathrm{iSt}(k,n)}T_{X}\mathrm{iSt}(k,n)$ to $\mathrm{iSt}(k,n)$ satisfying the following conditions for all $X\in\mathrm{iSt}(k,n):$
@@
where $\mathcal{R}_{X}$ is the restriction of $\mathcal{R}$ to $T_{X}\mathrm{iSt}(k,n).$ The crucial role of a retraction in optimization is to maintain the feasibility of the updated iterate from a feasible search direction which is a tangent vector at the current iterate. In TiepSon2025 , a retraction based on the Cayley transformation was proposed. In this work, based on the formulation of a quasi-geodesic (33), we propose another retraction which will be termed quasi-geodesic retraction as follows:
@@
Proof We only need to verify the conditions (R.1) and (R.2) since the expression in (41) is well-defined for any $X\in\mathrm{iSt}(k,n)$ and $Z\in T_{X}\mathrm{iSt}(k,n)$ thanks to Proposition 5.
@@
where we have invoked the identity $\frac{\mathrm{d}t}{t}\mathrm{e}^{M(t)}|_{t=0}=\mathrm{e}^{M(0)}\dot{M}(0)$ for smooth one-parameter matrix $M(t)$ such that $M(0)\dot{M}(0)=\dot{M}(0)M(0)$, see (Wilcox1967, , Eq. (2.1)). ∎
@@ 6 Numerical examples
In this section, theoretical results are assembled to enable a Riemannian gradient descent scheme and then numerically verified via examples.
@@ A Riemannian gradient descent method
To minimize a continuously differentiable cost function $f$ on the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}$,
@@
given an initial point $X_{0}\in\mathrm{iSt}_{A,J}(k,n)$, one consecutively updates the approximate solution by
@@
where $\mathcal{R}_{X_{j}}(\cdot)$ is a retraction at $X_{j}$, $\mathrm{grad}f(X_{j})$ denotes the Riemannian gradient of the cost function $f$ at $X_{j}$ with respect to a chosen Riemannian metric $\mathrm{g}(\cdot,\cdot)$, and $\tau_{j}>0$ is a step size to be determined. To compute the step size, we combine a nonmonotone line search technique with the alternating Barzilai-Borwein strategy on an embedded submanifold, see BarB88 ; HuLWY2020 ; IannP18 ; ZhanH2004 . For the sake of convenience, we recall the details in Algorithm 1. Note from (43) that if $\alpha$ is chosen to be zero, this scheme reduces to the popular monotonically decreasing line search with Amijo’s rule. Algorithm 1 was also employed in TiepSon2025 for numerical verification of optimization using the geometric quantities under a general tractable metric and the Cayley retraction. Its global convergence is guaranteed by (GSAS21, , Thm. 5.7). Namely, regardless of the initial point $X_{0}$, each accumulation point $X^{*}$ of the iterate generated by Algorithm 1 satisfies the first-order necessary optimality condition $\mathrm{grad}f(X^{*})=0$.
@@
For the backtracking search, we set $\beta=1\mathrm{e}{-4}$, $\delta=5\mathrm{e}{-1}$, $\gamma_{0}=1\mathrm{e}{-3}$, $\gamma_{\min}=1\mathrm{e}{-15}$, $\gamma_{\max}=1\mathrm{e}{+5}$. Moreover, $\alpha=0.85$ is used in the nonmonotone condition. A running is considered convergent when the norm of the Riemannian gradient of the cost function at the current iterate is relatively small compared to that at the initial guess $X_{0}$, namely, $\|\mathrm{grad}f(X_{j})\|_{X_{j}}\leq\texttt{rstop}\|\mathrm{grad}f(X_{0})\|_{X_{0}}$. In all tests, rstop is set to be $1\mathrm{e}-5$.
@@
where $n=1000,k=200$, $M=VV^{T}$ with $V$ is the $n\times(n-5)$ matrix randomly generated with default setting and then orthogonalized by orth in MATLAB. For the manifold, let $p=750,m=250,k_{p}=k_{m}=100$, we set $A=\mathrm{diag}(1,\ldots,p,-1,\ldots,-m)$ and $J=\mathrm{diag}(I_{k_{p}},-I_{k_{m}})$. The initial guess $X_{0}$ is set as the $n\times k$ zero matrix except for the submatrices $X_{0}(1,\ldots,k_{p};1,\ldots,k_{p})=\mathrm{diag}(\frac{1}{\sqrt{1}},\ldots,\frac{1}{\sqrt{k_{p}}})$ and $X_{0}(p+1,\ldots,p+k_{m};k_{p}+1,\ldots,k_{p}+k_{m})=\mathrm{diag}(\frac{1}{\sqrt{1}},\ldots,\frac{1}{\sqrt{k_{m}}})$.
@@
on the $J$-orthogonal group with $n=500$, $J=\mathrm{diag}(I_{p},-I_{m}),$ where $p=375,$ $m=125$. Numerical data are generated as follow: $G$ is a $(n-20)\times n$ random matrix with default setting and $B$ is set as $G\mathrm{diag}(U,V)$ where $U\in\mathbb{R}^{p\times p}$ and $V\in\mathbb{R}^{m\times m}$ are two orthogonalized random matrices generated with the setting rng (1,’twister’). The $n\times n$ identity matrix is used as the initial guess.
@@
It is worth to note that, while the Riemannian gradient is defined upon the metric, the retraction is metric-independent. Therefore, in Algorithm 1, one can combine any metric with any available retraction. Namely, there are the Cayley retraction and the one based on the quasi-geodesic developed in this paper. Regarding metrics, we have the general tractable and the generalized canonical metrics. In TiepSon2025 , the Euclidean and weighted Euclidean metrices have been used. The examples there showed that using the latter is much more advantageous. However, the Euclidean Hessian of the cost functions considered here are both degenerate; the choice of weighted Euclidean metric as in TiepSon2025 is inappropriate. Thus, we will use in our tests the Euclidean and the generalized canonical metrics. Consequently, there are totally four combinations whose corresponding numerical results will be marked by “Eucl-Cayley”, “Eucl-qgeo”, “gcan-Cayley”, and “gcan-qgeo”. Besides, since in our code we do not attempt to exploit the special structure of $A$, we always employ (17) for $X_{\perp}$. All the tests are done on a laptop with Intel(R) Core(TM) i7-4500U (at 1.8GHz, 5MB Cache, 8GB RAM) running MATLAB R2023a under Windows 10 Home. (The main code is made available at https://sites.google.com/view/ntson/code.)
We report the numerical results in Table 1 and Table 2 in which obj., grad., feas., #\#iter., #\#eval., and CPU represent the value of the cost function, the norm in the chosen metric of the Riemannian gradient of the cost function, the feasibility $\|X_{*}^{T}AX_{*}-J\|_{F}$, the number of iterations, the number of function evaluations, and the elapsed time, respectively. In the case that the iteration has not reached the stopping criterion, $X^{*}$ is replaced by the last iterate. In addition, the convergence histories are displayed in Figure 1 and Figure 2.
Table 1: Results for minimization of trace cost function (44).
@@
Eucl-Cayley $0.6387$ $4.0204\mathrm{e}-6$ $2.3354\mathrm{e}-13$ 389 423 97.3941
Eucl-qgeo $0.6387$ $3.9532\mathrm{e}-6$ $5.0332\mathrm{e}-08$ 392 437 50.3922
gcan-Cayley 0.6387 $4.2671\mathrm{e}-6$ $2.5933\mathrm{e}-13$ 472 516 93.2189
gcan-qgeo 0.6387 $4.1930\mathrm{e}-6$ $3.1692\mathrm{e}-13$ 458 598 47.1033
Table 2: Results for minimization of Procrustes cost function (45).
@@
Eucl-Cayley $21.5943$ $5.9039$ $2.5425\mathrm{e}-12$ 2000 2042 $2068$
Eucl-qgeo $38.1905$ $5.6773$ $8.4230\mathrm{e}-10$ 2000 2064 $2817$
gcan-Cayley 2.8671 $1.4755$ $3.3286\mathrm{e}-13$ 340 615 54
gcan-qgeo 15.5090 $1.5283$ $1.2650\mathrm{e}-10$ 402 738 356
@@ Discussion
Through examples presented, we can first see that, all newly established theoretical materials work properly and compatibly with the existing ones. Second, as expected, for the case that $k$ is not small, the schemes using the generalized canonical metric always need less time to be convergent (with respect to the given rstop) compared to the ones using the Euclidean metric; even in the case of Procrustes problem, within at most 2000 iterations, only the schemes with the generalized canonical metric converge. Third, the quasi-geodesic retraction apparently is less feasible than the Cayley retraction. This fact might come from the instability of the scaling and squaring algorithm invoked by expm in MATLAB which occasionally occurs as pointed out in GuetN2016 .

### v2

@@ 5 Quasi-geodesic based retraction on $\mathrm{iSt}(k,n)$
Let us first formally define a retraction on $\mathrm{iSt}(k,n)$; see AbsiMS08 ; AdleDMMS02 for more details about retractions on a general Riemannian manifold. It is a smooth mapping $\mathcal{R}$ from the tangent bundle $T\mathrm{iSt}(k,n)\coloneqq\bigcup_{X\in\mathrm{iSt}(k,n)}T_{X}\mathrm{iSt}(k,n)$ to $\mathrm{iSt}(k,n)$ satisfying the following conditions for all $X\in\mathrm{iSt}(k,n):$
@@
where $\mathcal{R}_{X}$ is the restriction of $\mathcal{R}$ to $T_{X}\mathrm{iSt}(k,n).$ The crucial role of a retraction in optimization is to maintain the feasibility of the updated iterate from a feasible search direction which is a tangent vector at the current iterate. In TiepSon2025 , a retraction based on the Cayley transformation was proposed. In this work, based on the formulation of a quasi-geodesic (32), we propose another retraction, which will be termed quasi-geodesic retraction, as follows:
@@
Proof We only need to verify the conditions (R.1) and (R.2) since the expression in (41) is well-defined for any $X\in\mathrm{iSt}(k,n)$ and $Z\in T_{X}\mathrm{iSt}(k,n)$ thanks to Proposition 5
@@
where we have invoked the identity $\frac{\mathrm{d}t}{t}\mathrm{e}^{M(t)}|_{t=0}=\mathrm{e}^{M(0)}\dot{M}(0)$ for smooth one-parameter matrix $M(t)$ such that $M(0)\dot{M}(0)=\dot{M}(0)M(0)$; see (Wilcox1967, , Eq. (2.1)). ∎
@@ 6 Numerical examples
In this section, several numerical examples are presented to verify our theoretical findings. As constructed in Section 3, two formulae for computing the Riemannian gradient using the proposed metric are given in (23) and (24); it was also stated that the latter outperforms the former when $A$ is dense. First, we validate this statement via an extra example. After that, the derived components are assembled to enable a Riemannian gradient descent scheme and then numerically verified via several model tests. The tests are done on a laptop with Intel(R) Core(TM) i7-4500U (at 1.8GHz, 5MB Cache, 8GB RAM) running MATLAB R2023a under Windows 10 Home. (The main code is made available at https://sites.google.com/view/ntson/code.)
A numerical comparison of (23) and (24)
We construct a manifold $\mathrm{iSt}_{A,J}$ as follows: the matrix $A=P\mathrm{diag}(1,2,\ldots,1500,-500,-499,\ldots,-1)P^{T}$, where $P$ is an orthogonalized random matrix and $J=\mathrm{diag}(I_{14},-I_{6})$. We take $f(X)=\frac{1}{2}\mathrm{tr}\left(X^{T}MX\right)$ with $M$ is a fixed spd matrix of apt size. We compute the Riemannian gradient using two formulae (23) and (24) at $X=\left[p_{1}/\sqrt{1},\ldots p_{14}/\sqrt{14},p_{1501}/\sqrt{500},\ldots,p_{1506}/\sqrt{495}\right]$ in which $p_{j}$ is the $j$-th column of $P$ generated previously. We run this setting 10 times and take the average time elapsed. The result shows that, given the Euclidean gradient $\nabla f(X)=MX$, while computing the Riemannian gradient following (23) takes about 0.45(s), the one invoking (24) needs only about 0.016(s), which is roughly 30 times faster. Therefore, only (24) is further used in our tests.
@@ A Riemannian gradient descent method
One option to minimize a continuously differentiable cost function $f$ on the indefinite Stiefel manifold $\mathrm{iSt}_{A,J}$,
@@
given an initial point $X_{0}\in\mathrm{iSt}_{A,J}(k,n)$, is to consecutively updates the approximate solution by
@@
where $\mathcal{R}_{X_{j}}(\cdot)$ is a retraction at $X_{j}$, $\mathrm{grad}f(X_{j})$ denotes the Riemannian gradient of the cost function $f$ at $X_{j}$ with respect to a chosen Riemannian metric $\mathrm{g}(\cdot,\cdot)$, and $\tau_{j}>0$ is a step size to be determined. To compute the step size, we combine a nonmonotone line search technique with the alternating Barzilai-Borwein strategy on an embedded submanifold; see BarB88 ; HuLWY2020 ; IannP18 ; ZhanH2004 . For the sake of convenience, we recall the details in Algorithm 1. Note from (43) that if $\alpha$ is chosen to be zero, this scheme reduces to the popular monotonically decreasing line search with Amijo’s rule. Algorithm 1 was also employed in TiepSon2025 for numerical verification of optimization using the geometric quantities under a general tractable metric and the Cayley retraction. Its global convergence is guaranteed by (GSAS21, , Thm. 5.7). Namely, regardless of the initial point $X_{0}$, each accumulation point $X^{*}$ of the iterate generated by Algorithm 1 satisfies the first-order necessary optimality condition $\mathrm{grad}f(X^{*})=0$.
@@
For the backtracking search, we set $\beta=1\mathrm{e}{-4}$, $\delta=5\mathrm{e}{-1}$, $\gamma_{0}=1\mathrm{e}{-3}$, $\gamma_{\min}=1\mathrm{e}{-15}$, $\gamma_{\max}=1\mathrm{e}{+5}$. Moreover, $\alpha=0.85$ is used in the nonmonotone condition. A running is considered convergent when the norm of the Riemannian gradient of the cost function at the current iterate is relatively small compared to that at the initial guess $X_{0}$. Namely, $\|\mathrm{grad}f(X_{j})\|_{X_{j}}\leq\texttt{rstop}\|\mathrm{grad}f(X_{0})\|_{X_{0}}$, where $\|\cdot\|_{X_{j}}$ is the norm associated with the chosen metric at $X_{j}$. In all tests, rstop is set to be $1\mathrm{e}-5$. We also stop the loop when $j$ exceeds a given maximal number of iterations iter which is set to $1000$.
@@
where $n=1000,k=500$, $M=VV^{T}$ with $V$ is the $n\times(n-5)$ matrix randomly generated with default setting and then orthogonalized using orth in MATLAB; matrix $M$ is finally normalized by dividing $M$ by its Frobenius norm. For the manifold, let $p=750,m=250,k_{p}=400,k_{m}=100$, we set $A=\mathrm{diag}(1,\ldots,p,-1,\ldots,-m)$ and $J=\mathrm{diag}(I_{k_{p}},-I_{k_{m}})$. The initial guess $X_{0}$ is, for the sake of simplicity, set as the $n\times k$ zero matrix except for the submatrices $X_{0}(1,\ldots,k_{p};1,\ldots,k_{p})=\mathrm{diag}(1/\sqrt{1},\ldots,1/\sqrt{k_{p}})$ and $X_{0}(p+1,\ldots,p+k_{m};k_{p}+1,\ldots,k_{p}+k_{m})=\mathrm{diag}(1/\sqrt{1},\ldots,1/\sqrt{k_{m}})$.
@@
on the $J$-orthogonal group High03 with $n=500$, $J=\mathrm{diag}(I_{p},-I_{m}),$ where $p=375,$ $m=125$. Numerical data are generated as follows: $G$ is a $(n-20)\times n$ random matrix with default setting and then normalized as above, and $B$ is set as $G\mathrm{diag}(U,V)$ where $U\in\mathbb{R}^{p\times p}$ and $V\in\mathbb{R}^{m\times m}$ are two orthogonalized random matrices generated with the setting rng (1,’twister’). The $n\times n$ identity matrix is used as the initial guess.
@@
It is worth noting that, while the Riemannian gradient is defined upon the metric, the retraction is metric-independent. Therefore, in Algorithm 1, one can combine any metric with any available retraction. Namely, there are the Cayley retraction and the one based on the quasi-geodesic developed in this paper. Regarding metrics, we have the general tractable and the generalized canonical metrics. In TiepSon2025 , the Euclidean and weighted Euclidean metrics have been used. The examples there showed that using the latter is much more advantageous. However, the Euclidean Hessian of the cost functions considered here are both degenerate; the choice of weighted Euclidean metric as in TiepSon2025 is inappropriate. Thus, we will use in our tests the Euclidean and the generalized canonical metrics. Consequently, there are totally four combinations whose corresponding numerical results will be marked by “Eucl-Cayley”, “Eucl-qgeo”, “gcan-Cayley”, and “gcan-qgeo”.
We report the numerical results in Table 1 and Table 2 in which obj., grad., feas., #\#iter., #\#eval., and CPU represent the value of the cost function, the norm in the chosen metric of the Riemannian gradient of the cost function, the feasibility $\|{X^{*}}^{T}AX^{*}-J\|_{F}$, the number of iterations, the number of function evaluations, and the elapsed time in second, respectively. In the case that the iteration has not reached the stopping criterion, $X^{*}$ is replaced by the last iterate. In addition, the convergence histories are displayed in Figure 1 and Figure 2 in which we present the value of the cost function, the norm of the Riemannian gradient, and the time elapsed along the iterations, respectively from the left to the right.
Table 1: Results for minimization of the trace cost function (44).
@@
Eucl-Cayley $0.0398$ $1.0383\mathrm{e}-7$ $5.1411\mathrm{e}-13$ 480 524 750
Eucl-qgeo $0.0398$ $1.2258\mathrm{e}-7$ $9.9156\mathrm{e}-08$ 407 446 685
gcan-Cayley 0.0398 $1.0342\mathrm{e}-7$ $6.6291\mathrm{e}-13$ 614 666 260
gcan-qgeo 0.0398 $1.1001\mathrm{e}-7$ $4.8909\mathrm{e}-13$ 461 505 324
Table 2: Results for minimization of the Procrustes cost function (45).
@@
Eucl-Cayley $9.0163\mathrm{e}-4$ $1.9212\mathrm{e}-3$ $1.0842\mathrm{e}-12$ 1000 1034 $1096$
Eucl-qgeo $7.6519\mathrm{e}-4$ $2.0001\mathrm{e}-4$ $8.4243\mathrm{e}-13$ 1000 1034 $1491$
gcan-Cayley $2.0092\mathrm{e}-4$ $2.0370\mathrm{e}-5$ $4.2354\mathrm{e}-13$ 435 757 70
gcan-qgeo $1.9463\mathrm{e}-4$ $1.8871\mathrm{e}-5$ $6.4464\mathrm{e}-13$ 402 734 400
@@ Discussion
Through examples presented, we can first see that, all newly established theoretical components work properly and compatibly with the existing ones. Second, as expected, for the cases that $k$ is not small in both models, the schemes using the generalized canonical metric need less time to be convergent with respect to a given rstop compared to the ones using the Euclidean metric although in some cases the required number of iterations can be larger. Third, the quasi-geodesic retraction apparently is less feasible than the Cayley retraction. This fact might come from the instability of the scaling and squaring algorithm invoked by expm in MATLAB which occasionally occurs as pointed out in GuetN2016 . And finally, we note that while our sole purpose in constructing the generalized canonical metric is to avoid solving the Lyapunov equation in computing the Riemannian gradient, altering the metric has a great effect on the aspects of numerical methods, such as (numerical) convergence, number of iterations, as one can see in our tests. The result can also depend on the model, namely the cost function. It is advised in MishS2016 that the Hessian of the cost function should be taken into account when constructing the metric. To understand the details, we believe that deeper and problem-dependent investigations are required.